# Mental Disorders Detection in the Era of Large Language Models

### v1

@@ 5.1.2 DSM Dataset
The best result achieved by Vikhr 7B IT 0.4 model, evaluated in 5-shot, with 66.1% F1-macro. The traditional machine learning methods, as well as encoder models, performed poorly with nearly 53% F1-macro on linguistic features.
Comparing the results between the DE and DSM datasets, a significant difference in classification quality can be observed. This may be due to several factors. First of all, we can note the sample size, which was significantly smaller in the DSM dataset. Anyway, almost all machine learning models, regardless of the selected features, can show low accuracy on a small amount of data. On the other hand, if we refer to the study (Ignatiev et al., 2022), where less stringent restrictions on text volume and temporal proximity to questionnaire screening dates were applied to the same raw data, the results were still not very high: about 60% F1-macro for psycholinguistic features and TF-IDF features.
Secondly, the problem may lie in the format of the texts from the social network. The texts in DSM are concatenated together with a collection of posts from users’ personal pages, and they mostly lack coherent logic and cohesion in the resulting text. Even with the 6,000 character limit, the standard deviation in word count in this data is on the order of 300 words, while in DE the similar value is around 100. Thus, DSM is a much noisier dataset, which can strongly affect the quality of classification with psycholinguistic markers, where the values of some markers can be affected by the volume of text analyzed, even considering their normalization with respect to the text volume.
Finally, the way in which the target class of pathology is defined can be of great importance. Although a widely used and validated psychological questionnaire was used for the DSM dataset, it is not possible to directly compare its results with a clinically confirmed diagnosis. The same findings can be outlined in work that criticizes the way in which social media users are labeled for mental illness by indirectly affiliating or self-identifying mental ill-health (Ernala et al., 2019). In favor of the significance of this factor is also the fact that TF-IDF-based features performed significantly better on DE data than on DSM, although, unlike psycholinguistic markers, they do not have an initial specialization for detecting signs of mental ill-health.
To take into account all of the aforementioned considerations, we applied LLMs to this task and showed, that the bigger model is able to partially overcome these issues. While RuBERT and the model trained on linguistic features both show comparatively low results, the evaluated in 5-shot format model performs significantly better. The same outcomes are observed in the experiments with anxiety datasets.
@@ 5.1.3 AL, AD, and AC Datasets
Turning to the anxiety detection task, the difference between the AutoML-based and encoder-based models is only noticeable on the picture description (AD) dataset, where the accuracy of the RuRoBERTa model was 55% F1-macro. Overall, it can be said that none of the non-LLM models performed well in this task. However, LLMs performed significantly better on these tasks. Models with LoRA in general perform better than encoder-based and AutoML-based models, but the best performance is achieved by 0-shot and 5-shot prompting. Again, this can be linked to the complexity of the domain and the small amount of training data, which makes it hard to finetune a model.
@@ 5.2 Classification on D-all and A-all Datasets
We combined the depression and anxiety datasets into pooled datasets D-all (DE and DSM) and A-all (AL, AD, and AC). Classification performance of the best models on D-all and A-all combined data presented in Table 3.
The LLMs performed better than other models on both D-all and A-all. It is also noticeable that methods with model tuning work better on D-all, while prompting shows better results on A-all. It can be explained with significant genre and length variations on A-all datasets, so the general-purpose models perform better than finetuned ones due to the complexity of the data.
@@ 5.3 Classification of DSM Data with Best Models from DE
The classification results of models that demonstrated the best performance on the DE dataset applied to DSM test data are shown in Table 4.
@@ 6 Conclusion
7 Limitations
The main limitation of this paper is that it is impossible to share all the raw texts from used datasets, as they are all distributed under different terms.
The amount of sample data used for predicting anxiety is small, which does not allow us to adequately judge the possibility of predicting anxiety in Russian text using machine learning methods. Although the results on all the anxiety datasets used show poor accuracy, perhaps a very different scale of data is needed for this task. The paper does not discuss the differences between the results inside one group of the models in detail as this was not the aim of the paper.
We conducted the experiments only for the Russian due to the poor availability of related datasets for other languages. However, the used methods in general are language agnostic so the results could be extended to other languages. The overall result about LLMs as the best-performing models matches similar studies for English on closed data.
8 Ethical Considerations
The problem discussed in this paper is the sensitive issue of mental health. To avoid any possible harm, we did not fully open-source used datasets. All shared data does not contain any information that names or uniquely identifies individuals, nor does it contain offensive content. The examined models for the detection of mental disorders do not aim to replace a professional physician, on the contrary, these models are intended to support a human expert.
@@ Appendix A Hyperparameters
The used net and optimal parameters, alongside with used checkpoint of encoder models, are presented in Table 6. The hyperparameters for LLMs, tuned with LoRA are presented in Table 7. The LoRA was applied to the following target layers: q_proj, up_proj, o_proj, k_proj, down_proj, gate_proj, v_proj. For LLMs 0-shot and 5-shot evaluation, we used the following generation parameters: do_sample=False, temperature=1.0, top_p=1.0, top_k=50, repetition_penalty=1.0. For MMLU-style evaluation generation length was limited to one token while for regular evaluation it was set to 64 tokens. Prompt examples are presented in Table 5.
@@ Appendix B Full Results for Models
Table 8 shows results for TF-IDF, models on linguistic features, and encoder-based transformers. Tables 9, 11 and 10 contain results for LLMs with PEFT, 0-shot and 5-shot evaluation, and 0-shot and 5-shot MMLU-style evaluation correspondingly.
@@
| DE | RuRoBERTa | 90.444.62 | 95.932.28 | 92.981.92 | 64.729.29 | 56.8225.81 | 60.3727.21 | 76.6714.52 | ||||||||
| RuBioRoBERTa | 91.925.21 | 95.193.05 | 93.362.13 | 64.2529.33 | 63.6428.63 | 63.7328.62 | 78.5415.3 | |||||||||
| RuBERT | 94.451.16 | 91.111.7 | 92.741.04 | 68.414.04 | 78.034.85 | 72.83.43 | 82.772.21 | |||||||||
| BERT | 92.970.8 | 92.961.23 | 92.960.85 | 71.343.96 | 71.213.39 | 71.233.14 | 82.091.98 | |||||||||
| Linguistic features | 94.00.8 | 95.21.0 | 94.60.8 | 79.34.0 | 75.03.5 | 77.03.3 | 85.82.1 | |||||||||
| TF-IDF | 91.62.1 | 94.42.2 | 93.01.2 | 74.87.0 | 64.410.3 | 68.56.6 | 80.73.8 | |||||||||
| DSM | RuRoBERTa | 61.314.38 | 73.4612.73 | 66.587.66 | 46.7610.29 | 31.486.93 | 36.686.2 | 51.635.8 | ||||||||
| RuBioRoBERTa | 62.12.98 | 62.966.05 | 62.464.3 | 43.664.43 | 42.594.14 | 43.013.7 | 52.743.67 | |||||||||
| RuBERT | 60.781.75 | 96.916.9 | 74.521.07 | 9.0920.33 | 5.5612.42 | 6.915.42 | 40.717.18 | |||||||||
| BERT | 61.091.9 | 69.1414.61 | 64.135.24 | 34.5915.71 | 33.3316.97 | 33.7116.01 | 48.925.62 | |||||||||
| Linguistic features | 62.81.8 | 59.37.4 | 60.74.0 | 43.72.7 | 47.27.7 | 45.13.8 | 52.92.0 | |||||||||
| TF-IDF | 62.22.6 | 51.213.9 | 55.38.5 | 42.93.6 | 53.711.4 | 46.94.7 | 51.13.5 | |||||||||
| AL | RuRoBERTa | 56.825.65 | 75.07.76 | 64.495.59 | 52.5410.76 | 33.3312.77 | 40.1712.48 | 52.338.31 | ||||||||
@@
| RuBERT | 53.633.2 | 77.2719.99 | 62.087.52 | 45.8329.56 | 21.9320.48 | 24.6517.57 | 43.365.79 | |||||||||
| BERT | 52.711.02 | 93.948.16 | 67.442.96 | 6.2513.98 | 2.635.88 | 3.78.28 | 35.572.99 | |||||||||
| Linguistic features | 47.421.4 | 36.417.8 | 40.919.2 | 48.52.7 | 68.414.6 | 56.13.7 | 48.58.2 | |||||||||
| TF-IDF | 61.55.0 | 46.23.1 | 52.62.3 | 51.22.5 | 65.87.9 | 57.54.5 | 55.02.9 | |||||||||
| AD | RuRoBERTa | 57.632.52 | 56.674.71 | 57.073.19 | 52.82.82 | 53.74.14 | 53.172.84 | 55.122.58 | ||||||||
| RuBioRoBERTa | 45.065.52 | 43.3319.72 | 41.4913.8 | 38.298.58 | 40.7423.28 | 37.7212.46 | 39.65.38 | |||||||||
| RuBERT | 48.637.83 | 80.029.58 | 59.3415.36 | 12.5918.14 | 12.9618.33 | 12.718.04 | 36.024.94 | |||||||||
| BERT | 52.031.0 | 93.337.99 | 66.742.88 | 20.5621.12 | 4.634.99 | 7.347.71 | 37.042.63 | |||||||||
| Linguistic features | 50.83.3 | 43.312.5 | 45.38.1 | 44.72.6 | 51.915.9 | 47.37.4 | 46.31.8 | |||||||||
| TF-IDF | 54.915.1 | 37.54.8 | 43.22.5 | 44.67.8 | 59.321.0 | 50.412.7 | 46.87.1 | |||||||||
| AC | RuRoBERTa | 45.720.48 | 77.0435.48 | 57.1925.66 | 24.5724.96 | 24.5635.49 | 21.2223.82 | 39.27.55 | ||||||||
| RuBioRoBERTa | 52.921.41 | 76.314.15 | 62.095.54 | 34.8815.72 | 20.1811.64 | 24.8412.88 | 43.474.21 | |||||||||
| RuBERT | 54.291.57 | 87.0419.35 | 66.08.24 | 43.5922.45 | 14.0417.1 | 17.1515.7 | 41.584.88 | |||||||||

### v3

@@ 5.1.2 DSM Dataset
The best result achieved by the Vikhr 7B IT 0.4 model, evaluated in the 5-shot regime, with 63.69% F1-macro. The traditional machine learning methods, as well as encoder models, performed poorly with nearly 53% F1-macro on linguistic features.
Comparing the results between the DE and DSM datasets, a significant difference in classification quality can be observed. This may be due to several factors. First of all, the sample size is significantly smaller in the DSM dataset, as machine learning models in general can show low accuracy on limited data. On the other hand, if we refer to the study (Ignatiev et al., 2022), where less stringent restrictions on text volume and temporal proximity to questionnaire screening dates were applied to the same raw data, the results were still not very high: about 60% F1-macro for psycholinguistic features and TF-IDF features.
Secondly, the text format in the DSM introduces some noise. The texts in DSM are concatenated together with a collection of posts from users’ personal pages, and they mostly lack coherent logic and cohesion in the resulting text. Even with the 6,000-character limit, the standard deviation in word count is approximately 300 words, compared with 100 in DE. Thus, DSM is a much noisier dataset, which can strongly affect the quality of classification with psycholinguistic markers, where the values of some markers can be affected by the volume of text analyzed, even considering their normalization with respect to the text volume.
Finally, the way in which the target class of pathology is defined can be of great importance. Although a widely used and validated psychological questionnaire was used for the DSM dataset, its results cannot be compared with a clinically confirmed diagnosis. The same findings can be outlined in work that criticizes the way in which social media users are labeled for mental illness by indirectly affiliating or self-identifying mental ill-health (Ernala et al., 2019). In favor of the significance of this factor is also the fact that TF-IDF-based features performed significantly better on DE data than on DSM, although, unlike psycholinguistic markers, they do not have an initial specialization for detecting signs of mental ill-health.
To take into account these considerations, we applied LLMs to this task and showed that the bigger model is able to partially overcome these issues. While RuBERT and the model trained on linguistic features both show comparatively low results, a 5-shot LLM performs significantly better. The same outcomes are observed in the experiments with anxiety datasets.
For a better understanding of per-class performance of models, we also provided results for the best models on the DSM dataset for various splits within a class based on the initial BDI scores. The results are provided in Appendix˜F.
@@ 5.1.3 AL, AD, and AC Datasets
Turning to the anxiety detection task, the difference between the AutoML-based and encoder-based models is only noticeable on the picture description (AD) dataset, where the accuracy of the RuRoBERTa model was 55% F1-macro. Overall, it can be said that none of the non-LLM models performed well in this task. However, LLMs performed significantly better on these tasks. Models with LoRA, in general, perform better than encoder-based and AutoML-based models, but the best performance is achieved by 0-shot and 5-shot prompting. Again, this can be linked to the complexity of the domain and the small amount of training data, which makes it hard to fine-tune a model.
@@ 5.2 Classification on D-all and A-all Datasets
We combined the depression and anxiety datasets into pooled datasets D-all (DE and DSM) and A-all (AL, AD, and AC). Classification performance of the best models on D-all and A-all combined data is presented in Table˜3.
The LLMs performed better than other models on both D-all and A-all. It is also noticeable that methods with model tuning work better on D-all, while prompting shows better results on A-all. It can be explained with significant genre and length variations in A-all datasets, so the general-purpose models perform better than finetuned ones due to the complexity of the data.
@@ 5.3 Classification of DSM Data with Best Models from DE
The classification results of models that demonstrated the best performance on the DE dataset applied to DSM test data are shown in Table˜4.
@@
5.4 LLM Results Interpretation
As shown in Section˜5.3, the overall best results were obtained on the DE dataset with LLMs. To further analyze the quality of these results from the clinical perspective, we conducted an additional experiment on LLM results interpretation.
We chose the best LLM in a 5-shot setting on the DE dataset (Vikhr 7B IT 5.4) and asked two psychologists with relevant experience to rate the LLM generation on a scale from 1 to 5. The LLM-generated answer in this setting consists of a predicted class label and a detailed explanation of why this label was chosen. Details of the rating scale and statistics are provided in Appendix˜D. The psychologists assessed only texts from patients with clinically diagnosed depression that the LLM correctly labeled.
The average score from expert psychologists for LLM explanations was assigned to 2.84 out of 5. Moreover, the psychologists also noted each explanation, which can be marked as true to some extent (e.g., with some true claims in the explanation) – there are approximately 66% of such explanations. These results show that the explanations for depression detection from texts contain both correct and erroneous parts. Overall, the LLM (in the described settings) does not generate enough explanations that meet the requirements of clinicians.
To further extend the error analysis in the LLM generations in the psychological domain, we asked psychologists to describe the most common errors in the generated explanations. To do so, psychologists annotated each explanation with the list of errors and categorized these mistakes into the following groups: (1) tautology, (2) groundless generalization, (3) false conclusion, (4) confabulation, (5) distortion of medical understanding of depression, (6) incompleteness of selected signs of depression. Description and examples of these error types are located in Table˜17 in the Appendix˜D. It should also be noted that from the general NLP perspective, most of these types of errors can be defined as hallucinations; however, we suppose that a more precise definition is needed in this specific domain.
On average, each explanation contains two or more errors. The most common types of errors are groundless generalization, false conclusion, and confabulation, which occur in 56%, 56%, and 50% of samples, respectively. However, these errors appear several times in each explanation. The incompleteness of selected signs of depression appears in 44%, while distortion of medical understanding of depression and tautology are the rare errors, which appear in 19% and 13%. These results show that there exist various types of errors, specific to the mental health disorders domain. The additional results with detailed analysis of feature importance for the pathology and healthy class prediction are located in Appendix˜E.
@@ 6 Conclusion
Finally, we evaluated LLM-generated explanations and showed that in the current state, these explanations do not meet the clinicians’ requirements with an overall score of 2.84 out of 5, which answers RQ3. We also constructed a detailed classification of common errors in LLM explanations from the clinicians’ perspective, to guide further improvements of LLMs in this domain.
Limitations
The main limitation of this paper is that it is impossible to share all the raw texts from the used datasets, as they are all distributed under different terms. However, we present several anonymized textual examples and extracted psycholinguistic features.
The amount of sample data used for predicting anxiety is small, which does not allow us to adequately judge the possibility of predicting anxiety in Russian text using machine learning methods. Although the results on all the anxiety datasets used show poor accuracy, perhaps a very different scale of data is needed for this task. The paper does not discuss the differences between the results within one group of the models in detail, as this was not the aim of the paper.
We conducted the experiments only for Russian due to the poor availability of related datasets for other languages. However, the used methods in general are language agnostic, so the results could be extended to other languages. The overall result about LLMs as the best-performing models matches similar studies for English on closed data.
The final limitation is the temporal validity of the results. Due to the fast growth of LLM-based solutions, future LLMs could outperform the obtained results. However, we suppose that the current state of the field already presents an interest and is therefore investigated in the paper.
Ethical Considerations
The problem discussed in this paper is the sensitive issue of mental health. To avoid any possible harm, we did not fully open-source the used datasets. All shared data does not contain any information that names or uniquely identifies individuals, nor does it contain offensive content. The examined models for the detection of mental disorders do not aim to replace a professional physician; on the contrary, these models are intended to support a human expert.
Acknowledgements
The work of Gleb Kuzmin was supported by a grant, provided by the MED of Russia (agreement identifier 000000C313925P4G0002) and the agreement with the ISP RAS dated June 20, 2025 No. 139-15-2025-011.
The research was carried out using the infrastructure of the shared research facilities «High Performance Computing and Big Data » of FRC CSC RAS (CKP «Informatics»)
@@ Appendix A Hyperparameters
The used net and optimal parameters, alongside the used checkpoint of encoder models, are presented in Table˜7. For the models trained on the psycholinguistic features and TF-IDF, the best model was selected with the AutoML pipeline, with the most models chosen as a Random Forest ensemble of up to four models.
The hyperparameters for LLMs, tuned with LoRA, are presented in Table˜8. The LoRA was applied to the following target layers: q_proj, up_proj, o_proj, k_proj, down_proj, gate_proj, v_proj. For LLMs 0-shot and 5-shot evaluation, we used the following generation parameters: do_sample=False, temperature=1.0, top_p=1.0, top_k=50, repetition_penalty=1.0. The predicted classes were extracted using the string matching algorithm. For MMLU-style evaluation, the generation length was limited to one token, while for regular evaluation, it was set to 64 tokens. Prompt examples are presented in Table˜5. For the 0-shot and 5-shot evaluation, we averaged the results on several prompts with various system parts to take into account the sensitivity of the generation with respect to the prompt. These prompt variations are presented in Table˜6. In all 0-shot and 5-shot experiments, we did not conduct best-of-N aggregation of answers and used one extracted answer per generation, which was averaged over several prompts to ensure generalizability of results.
The used hardware, as well as GPU hours, carbon footprint, and memory requirements are presented in Table˜9.
@@
| Task | Prompt number | Prompt example (only system) |
| Depression | 1 | System: Read the provided text and determine whether the author has signs of depression. Use the scale: 0 - no depression, 1 - depression. Then explain why you chose this option. |
| Depression | 2 | System: Evaluate the text for the author’s depressive state. Scale: 0 - no depression, 1 - depression is present. Justify your choice after indicating the answer. |
| Depression | 3 | System: Analyze this text and identify the presence of depressive manifestations in its author. Use a binary assessment: 0 - no depression, 1 - presence of depression. Provide a rationale for your decision. |
| Depression | 4 | System: Assess the psychological state of the author of the text in the context of depression. Use the gradation: 0 - no signs of depression, 1 - there are signs of depression. Detail the reasons for making your decision. |
| Anxiety | 1 | System: Read the provided text and determine whether the author has signs of anxiety. Use the scale: 0 - no anxiety, 1 - anxiety. Then explain why you chose this option. |
| Anxiety | 2 | System: Evaluate the text for the author’s anxiety. Scale: 0 - no anxiety, 1 - anxiety is present. Justify your choice after indicating the answer. |
| Anxiety | 3 | System: Analyze this text and identify the presence of anxiety in its author. Use a binary assessment: 0 - no anxiety, 1 - anxiety present. Justify your decision. |
| Anxiety | 4 | System: Assess the psychological state of the author of the text in the context of anxiety. Use the gradation: 0 - no signs of anxiety, 1 - signs of anxiety present. Detail the reasons for making your decision. |
@@
| GPU type | NVIDIA V100 32GB | NVIDIA A100 80 GB | NVIDIA H100 80 GB | Total |
| GPU Hours | 119 | 371 | 43 | 533 |
| Carbon footprint, kg CO2 | 11.50 | 29.87 | 3.46 | 44.83 |
@@ Appendix B Full Results for Models
Table˜10 shows results for TF-IDF, models on linguistic features, and encoder-based transformers. Tables˜11, 13 and 12 contain results for LLMs with PEFT, 0-shot and 5-shot evaluation, and 0-shot and 5-shot MMLU-style evaluation correspondingly.
@@
| DE | RuRoBERTa | 90.44 4.62 | 95.93 2.28 | 92.98 1.92 | 64.70 29.29 | 56.82 25.81 | 60.37 27.21 | 76.67 14.52 | ||||||||
| RuBioRoBERTa | 91.92 5.21 | 95.19 3.05 | 93.36 2.13 | 64.25 29.33 | 63.64 28.63 | 63.73 28.62 | 78.54 15.30 | |||||||||
| RuBERT | 94.45 1.16 | 91.11 1.70 | 92.74 1.04 | 68.41 4.04 | 78.03 4.85 | 72.80 3.43 | 82.77 2.21 | |||||||||
| BERT | 92.97 0.80 | 92.96 1.23 | 92.96 0.85 | 71.34 3.96 | 71.21 3.39 | 71.23 3.14 | 82.09 1.98 | |||||||||
| Linguistic features | 94.00 0.80 | 95.20 1.00 | 94.60 0.80 | 79.30 4.00 | 75.00 3.50 | 77.00 3.30 | 85.80 2.10 | |||||||||
| TF-IDF | 91.60 2.10 | 94.40 2.20 | 93.00 1.20 | 74.80 7.00 | 64.40 10.30 | 68.50 6.60 | 80.70 3.80 | |||||||||
| DSM | RuRoBERTa | 61.31 4.38 | 73.46 12.73 | 66.58 7.66 | 46.76 10.29 | 31.48 6.93 | 36.68 6.20 | 51.63 5.80 | ||||||||
| RuBioRoBERTa | 62.10 2.98 | 62.96 6.05 | 62.46 4.30 | 43.66 4.43 | 42.59 4.14 | 43.01 3.70 | 52.74 3.67 | |||||||||
| RuBERT | 60.78 1.75 | 96.91 6.90 | 74.52 1.07 | 9.09 20.33 | 5.56 12.42 | 6.90 15.42 | 40.71 7.18 | |||||||||
| BERT | 61.09 1.90 | 69.14 14.61 | 64.13 5.24 | 34.59 15.71 | 33.33 16.97 | 33.71 16.01 | 48.92 5.62 | |||||||||
| Linguistic features | 62.80 1.80 | 59.30 7.40 | 60.70 4.00 | 43.70 2.70 | 47.20 7.70 | 45.10 3.80 | 52.90 2.00 | |||||||||
| TF-IDF | 62.20 2.60 | 51.20 13.90 | 55.30 8.50 | 42.90 3.60 | 53.70 11.40 | 46.90 4.70 | 51.10 3.50 | |||||||||
| AL | RuRoBERTa | 56.82 5.65 | 75.00 7.76 | 64.49 5.59 | 52.54 10.76 | 33.33 12.77 | 40.17 12.48 | 52.33 8.31 | ||||||||
@@
| RuBERT | 53.63 3.20 | 77.27 19.99 | 62.08 7.52 | 45.83 29.56 | 21.93 20.48 | 24.65 17.57 | 43.36 5.79 | |||||||||
| BERT | 52.71 1.02 | 93.94 8.16 | 67.44 2.96 | 6.25 13.98 | 2.63 5.88 | 3.70 8.28 | 35.57 2.99 | |||||||||
| Linguistic features | 47.40 21.40 | 36.40 17.80 | 40.90 19.20 | 48.50 2.70 | 68.40 14.60 | 56.10 3.70 | 48.50 8.20 | |||||||||
| TF-IDF | 61.50 5.00 | 46.20 3.10 | 52.60 2.30 | 51.20 2.50 | 65.80 7.90 | 57.50 4.50 | 55.00 2.90 | |||||||||
| AD | RuRoBERTa | 57.63 2.52 | 56.67 4.71 | 57.07 3.19 | 52.80 2.82 | 53.70 4.14 | 53.17 2.84 | 55.12 2.58 | ||||||||
| RuBioRoBERTa | 45.06 5.52 | 43.33 19.72 | 41.49 13.80 | 38.29 8.58 | 40.74 23.28 | 37.72 12.46 | 39.60 5.38 | |||||||||
| RuBERT | 48.63 7.83 | 80.00 29.58 | 59.34 15.36 | 12.59 18.14 | 12.96 18.33 | 12.70 18.04 | 36.02 4.94 | |||||||||
| BERT | 52.03 1.00 | 93.33 7.99 | 66.74 2.88 | 20.56 21.12 | 4.63 4.99 | 7.34 7.71 | 37.04 2.63 | |||||||||
| Linguistic features | 50.80 3.30 | 43.30 12.50 | 45.30 8.10 | 44.70 2.60 | 51.90 15.90 | 47.30 7.40 | 46.30 1.80 | |||||||||
| TF-IDF | 54.90 15.10 | 37.50 4.80 | 43.20 2.50 | 44.60 7.80 | 59.30 21.00 | 50.40 12.70 | 46.80 7.10 | |||||||||
| AC | RuRoBERTa | 45.70 20.48 | 77.04 35.48 | 57.19 25.66 | 24.57 24.96 | 24.56 35.49 | 21.22 23.82 | 39.20 7.55 | ||||||||
| RuBioRoBERTa | 52.92 1.41 | 76.30 14.15 | 62.09 5.54 | 34.88 15.72 | 20.18 11.64 | 24.84 12.88 | 43.47 4.21 | |||||||||
| RuBERT | 54.29 1.57 | 87.04 19.35 | 66.00 8.24 | 43.59 22.45 | 14.04 17.10 | 17.15 15.70 | 41.58 4.88 | |||||||||